OpenStudy (anonymous):

use rational root theorum and the factor theorem to help solve the following equation x^3+x^2-25x-25

OpenStudy (mertsj):

Use synthetic to test the possible rational roots. You will find that -1 works and so x+1 is a factor and x^2+25 is the other factor.

OpenStudy (ash2326):

Using the rational root theorem we can guess the roots of an equation The rational roots maybe the factors of constant term over the coefficient of the highest power of x so here the roots maybe \[\frac{\pm 5 , 5,1}{1}\] Let's check if +5 is a root \[x^3+x^2-25x-25\] x=5 \[5^3+5^2-25 \times 5-25=125+25-125-25=0\] so 5 is a root or x-5 is a factor We have now \[x^3+x^2-25x-25\] \[x^2(x-5)+6x(x-5)+5(x-5)\]so we have \[(x^2+6x+5)(x-5)\] In this quadratic x^2+6x+5 the roots maybe given as \[\frac{\pm 5,1}{1}\] Let's check if -5 is a root \[x^2+6x+5=25-30+5=0\] so -5 is a root or x+5 is a factor Let's check -1 is a root, x=-1 \[1-6+5=0\] so (x+1) is also a root the roots are -1, 5 and -5